SM5 problems      (last update 06.10.05)

Xtal/TT Problem Solution
TT41 Always has 10-25% of bad events for run type PEDESTAL: the word count is correct, but the content of the data is not. TEST PULSE and LASER are insentive to this problem. We still don't know what is the reason - bad link or bad data from the FE card. 29-09-05: Solution found increasing the I2C setting to for laser-diode bias current for the GOH to 32. It was in default = 12. The maximum is 128. The solution applies to run 10271 and beyond.
general Very often 1 or 2 TT have a few % of bad events, but these towers are not always 52 and 57, but more or less random.
This effect is not understood yet also. During certain period, ~ 2 days, TT52 had also 20% of data integrity errors as TT41.
The same solution as avobe.
220/12 Has no data for the three gains, although the gain value it there.
It was working in 867, but never worked in H4 (always has zeros in data).
05-10-05: Partial solution found
803/33 Noisy
803 died about 1 month ago, most probably APD is broken. Paolo Rumerio reported all the story of this channel: (talk)
 
656/28 656 and other several channels are noisy. Much more information you can find in monitoring data base.  
TT 62 02-10-05: 3-6% of integrity errors in run H4-000012302-SM5-PEDESTAL-STD  
TT18 02-10-05: < 1 % of integrity errors in run H4-000012280-SM5-TEST_PULSE-MGPA  
TT43 02-10-05: 15-20% of integrity errors in run H4-000012272-SM5-TEST_PULSE-MGPA  
TT 8, 21, 63 02-10-05:  resp.: 5-20%, 40-40%, < 1% of integrity errors in run H4-000012192-SM5-TEST_PULSE-MGPA  
TT30 02-10-05: < 1 % of integrity errors in runs H4-000012178-SM5-LASER-STD and H4-000012180-SM5-LASER-STD  
TT13 02-10-05: 1-7 % of integrity errors in runs H4-000012167-SM5-PEDESTAL-STD  
TT 34 (40-60%) 02-10-05: 40-60 % of integrity errors in runs:

H4-000012148-SM5-PEDESTAL-STD
H4-000012150-SM5-TEST_PULSE-MGPA
H4-000012154-SM5-TEST_PULSE-MGPA
H4-000012162-SM5-TEST_PULSE-MGPA
H4-000012166-SM5-TEST_PULSE-MGPA

 
TT 41 02-10-05: 10-20 % of integrity errors in run H4-000012134-SM5-LASER-STD  
TT49 02-10-05: 15-20 % errors in:

LASER H4 11678 STD mon_done tape done 2005-10-02 18:58:22 CI LS
PEDESTALH4 11677 STD mon_done tape done 2005-10-02 18:57:32 CI PD
COSMIC H4 11676 STD mon_done tape done 2005-10-02 18:26:32 *CI PD CS
COSMIC H4 11675 STD mon_done tape done 2005-10-02 18:21:15 *CI PD CS
COSMIC H4 11674 STD mon_done tape done 2005-10-02 18:16:13 *CI PD CS
COSMIC H4 11673 STD mon_done tape done 2005-10-02 18:10:56 *CI PD CS
COSMIC H4 11672 STD mon_done tape done 2005-10-02 18:05:37 *CI PD CS
COSMIC H4 11671 STD mon_done tape done 2005-10-02 18:00:23 *CI PD CS
COSMIC H4 11670 STD mon_done tape done 2005-10-02 17:55:10 *CI PD CS
COSMIC H4 11669 STD mon_done tape done 2005-10-02 17:49:51 *CI PD CS TEST_PULSE H4 11668 MGPA mon_done tape done 2005-10-02 17:49:33 CI PD MP

 
1530/62 01-10-05: laser runs showing problems blue laser and red laser  
1672/67 01-10-05: laser runs showing problems blue laser (for any offset value the pedestal stays around 2300 ADC counts).  
1255/51 01-10-05: laser runs showing problems red laser